OpenStudy (anonymous):

Find the slope of the given line. 8y = 2x + 4

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

first convert to y=mx+b form, then m is your slope. So y = (2x+4)/8 y=1/4 * x + 1/2 1/4 is your slope :)